OpenStudy (anonymous):

A student used 10 mL water instead of 30mL for the extraction of the NaCl from the mixture. How would this affect the calculated percentage of NaCl in the sample?

OpenStudy (anonymous):

im confused a bit, what is he doing? taking salt and then dissolving it in 10 ml instead of 30 ml?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

if he dissolves same amount of salt in 10 ml and in 30 ml he would get higher percentage of salt in 10 ml and lower percentage in 30 ml

OpenStudy (anonymous):

Can u explain how? Well how you came up with that answer please?

OpenStudy (anonymous):

Just to better understand.

OpenStudy (anonymous):

lets say that he dissolves 5 g of anything in 10 and 30 ml lets say we calculate mass % gama(mass percentage) = m/V = 5g/10ml = 0,5 g ml-1 gama = 5g/30 ml=0,16 g ml-1

OpenStudy (anonymous):

Thanks alot.